# Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series

### v2

@@ 4 Theoretical properties
In this section, we investigate the optimization and statistical properties of TWI. We first show Algorithm 1 converges to critical points. Then, asymptotic consistency of TWI is established when the missing data concentrate on one side of $n_{1}$. For general non-systematic missing patterns, we show that identification of the underlying marginal distributions can be achieved by TWI under the assumptions of interpolator stability using a two-state Markov chain example (see (C1) and (C2) in Section 4.2).
Before presenting the results, it is worthwhile to clarify the asymptotic regime under consideration. The analysis is meaningful only when the number of missing entries is non-negligible compared to the sample size. Otherwise, if $\sharp(\mathcal{M})=o(n)$, where $\sharp(E)$ is the cardinality of the set $E$, then the empirical marginal distributions of any (bounded) imputations satisfy
$\frac{1}{n}\sum_{t\in\mathcal{A}_{n}}\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}+\frac{1}{n}\sum_{t\notin\mathcal{A}_{n}}\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}\Rightarrow\mu_{p},$
where $\mathcal{A}_{n}=\{t\in\{p-1,\ldots,n-1\}:\{t,t-1,\ldots,t-p+1\}\cap\mathcal{M}\neq\emptyset\}$. Therefore, assuming $\sharp(\mathcal{M})=o(n)$ would essentially preclude distributional inconsistency in the downstream. To avoid this trivial case, in this paper we maintain instead the assumption that $\mathcal{M}=\mathcal{M}_{n}$ is a sequence of index sets (independent of the underlying time series) such that $\sharp(\mathcal{M}_{n})/n\asymp 1$, where $a_{n}\asymp b_{n}$ means there exists $0<\underline{c}<\bar{c}<\infty$ such that $\underline{c}\leq a_{n}/b_{n}\leq\bar{c}$ for all large $n$.
@@ 4.1 Optimization and statistical guarantees
We begin by studying the optimization properties of Algorithm 1. All proofs are relegated to Section S1 of the supplementary material.
@@
    every limit point $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ of $\{(\hat{\mathbf{w}}^{(t+1)},\hat{\mathbf{\Pi}}^{(t)}):t=1,2,\ldots\}$ is a critical point of $F$. i.e.,
@@
If, in addition, the cost function $L$ takes the form $L(\mathbf{u},\mathbf{v})=\sum_{k=1}^{p}\ell(u_{k}-v_{k})$, where $\mathbf{u}=(u_{1},\ldots,u_{p})^{\top}$, $\mathbf{v}=(v_{1},\ldots,v_{p})^{\top}$, and $\ell$ is twice differentiable and convex, then
@@
    $F(\mathbf{w},\mathbf{\Pi})$ is biconvex in $\mathbf{w}$ and $\mathbf{\Pi}$.
Due to the biconvexity of $F$, one can employ off-the-shelf solvers for the subproblems in Algorithm 1. Moreover, it follows directly from the proof that if $\lambda>0$, $F(\mathbf{w},\mathbf{\Pi})$ is $\lambda$-strongly convex in $\mathbf{w}$. Hence, step (b) in Algorithm 1 admits a unique solution in each iteration and can be solved efficiently.
A few remarks are in order. First, it is straightforward to generalize Proposition 3 for multivariate time series. Second, like most methods involving non-convex optimization, a good initialization is crucial for converging to a “good” critical point. In our simulation studies, we use simple linear interpolation or ARIMA-based Kalman smoothing as initializations and TWI is able to find an imputation that offers substantial improvements over these initializations. Third, unlike many non-convex optimization problems, we are not satisfied by an arbitrary global minimizer of $F$. Essentially, it is because among all processes that have stable $p$-dimensional marginal distributions, only a subset of them matches the underlying marginal distributions. In Section 4.2, we probe this identification issue in more detail. However, when the missing data happen to lie on the same side of $t=n_{1}$, the global minimizer of $F$ achieves asymptotic consistency in the sense of Proposition 4 below.
@@
Assume (A1) and (A2) in Section S1 of the supplementary material. Suppose for any $\mathbf{w}\in\mathcal{C}$, $w_{t}=\tilde{x}_{t}$ if $t\notin\mathcal{M}_{n}$. Suppose further $n$ and $n_{1}$ satisfy $n_{1}\rightarrow\infty$ and $n-n_{1}\rightarrow\infty$. If either $\mathcal{M}_{n}\subseteq\{0,1,\ldots,n_{1}\}$ or $\mathcal{M}_{n}\subseteq\{n_{1}+1,\ldots,n-1\}$ (but not both), then, the global minimizer $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ of $F$ with $\lambda=0$ and $L(\mathbf{u},\mathbf{v})=\|\mathbf{u}-\mathbf{v}\|^{k}$, subject to $\mathbf{w}\in\mathcal{C}$ and the coupling conditions, satisfies
@@ 4.2 General missing pattern: A two-state Markov chain example
In this subsection, the identification issue of the temporal Wasserstein imputation method is studied via a two-state Markov chain example. Particularly, we connect the non-identification of the underlying marginal distributions to the lack of uniqueness of the solutions to a system of linear equations. We also show that, under two natural conditions, the correct marginal distributions can still be pinned down by TWI. To free our discussion from finite-sample complications, we will focus on the distribution level in this subsection, or intuitively speaking, assume “infinite sample.”
@@
Suppose the underlying time series $\{\tilde{x}_{t}\}_{-\infty<t<\infty}$ under study is a two-state Markov chain. Specifically, $\tilde{x}_{t}\in\{0,1\}$ and the transition probabilities are given by
$\mathbb{P}(\tilde{x}_{t}=1|\tilde{x}_{t-1}=0)=0.6,\quad$ $\mathbb{P}(\tilde{x}_{t}=0|\tilde{x}_{t-1}=1)=0.2.$
The transition probabilities are specified to facilitate discussions, but the main results below are not specific to these choices. It is easy to check that $\{\tilde{x}_{t}\}$ has the stationary distribution given by $\mathbb{P}(\tilde{x}_{t}=1)=0.75$ and $\mathbb{P}(\tilde{x}_{t}=0)=0.25$. Consequently, the 2-dimensional marginal distribution of $\{\tilde{x}_{t}\}$ is specified by
$\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(1,1))=0.6,\quad\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(1,0))=0.15,$ (9)
$\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(0,1))=0.15,\quad\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(0,0))=0.1.$
@@
Without loss of generality, we set $n_{1}=0$. Suppose that before time $n_{1}$, there is a missing value every $q_{1}$ observations. Likewise, there is a missing value every $q_{2}$ observations after time $n_{1}$. Specifically, $x_{t}=\bigstar$ if $-t$ is a positive integer multiple of $q_{1}$, and similarly, $x_{t}=\bigstar$ if $t$ is a positive integer multiple of $q_{2}$, where $q_{1}$ and $q_{2}$ are some integers larger than 2 and $q_{1}\neq q_{2}$ (see condition (C1) below). Next, suppose an imputation $\{\hat{w}_{t}\}_{-\infty<t<\infty}$ has been made such that if $t<n_{1}$,
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1)|(x_{t},x_{t-1})=(\bigstar,1))=a_{1},$ (10)
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0)|(x_{t},x_{t-1})=(\bigstar,0))=b_{1},$
and if $t\geq n_{1}$,
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1)|(x_{t},x_{t-1})=(\bigstar,1))=a_{2},$ (11)
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0)|(x_{t},x_{t-1})=(\bigstar,0))=b_{2},$
for some $a_{1},a_{2},b_{1},b_{2}\in[0,1]$. Note that $\mathbb{P}$ in (10) and (11) should be interpreted as proportions since the imputation may depend on the whole sequence of $\{x_{t}\}$, and it is assumed that they are well-defined. Essentially, here we parametrize all imputation methods by $a_{1},a_{2},b_{1},b_{2}$, which correspond to different 2-dimensional marginal distributions. Indeed, from (9)–(11), we can compute the 2-dimensional marginal distributions of the imputed series. For $\{\hat{w}_{t}\}_{-\infty<t<0}$,
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1))$ $=0.6(1-\frac{1}{q_{1}})+\frac{3a_{1}}{4q_{1}},\quad\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0))=0.15(1-\frac{1}{q_{1}})+\frac{b_{1}}{4q_{1}},$
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(0,1))$ $=0.15(1-q_{1}^{-1})+\frac{3}{4q_{1}}(1-a_{1}),$
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(0,0))$ $=0.1(1-q_{1}^{-1})+\frac{1}{4q_{1}}(1-b_{1}).$
@@
For $\{\hat{w}_{t}\}_{0\leq t<\infty}$, we can obtain similar expressions with $a_{1},b_{1},q_{1}$ replaced by $a_{2},b_{2},q_{2}$. The “correct” imputation should calibrate $a_{i}$’s and $b_{i}$’s such that the resulting 2-dimensional marginal distributions equal the marginal distribution of the underlying series, given in (9). To achieve such identification, the correct imputation must choose
$a_{1}^{*}=a_{2}^{*}=0.8,\quad b_{1}^{*}=b_{2}^{*}=0.6.$ (12)
@@
Note that this calibration does not depend on $q_{1},q_{2}$; it only depends on the distributional properties of the underlying process.
@@
Assuming the optimization is carried out perfectly [^1] This is where we ignore finite sample complications and assume a zero Wasserstein loss is achieved. One may use, for example, the hamming distance $L(x,y)=\mathbf{1}\{x\neq y\}$ as the ground metric for the discrete data., the temporal Wasserstein imputation (with $p\geq 2$) calibrates $a_{i},b_{i},i=1,2$, by equating the 2-dimensional marginal distributions of $\{\hat{w}_{t}\}_{-\infty<t<0}$ and $\{\hat{w}_{t}\}_{0\geq t<\infty}$. Hence, $\{a_{1},a_{2},b_{1},b_{2}\}$ satisfies
[…]
If $q_{1}=q_{2}$, then any imputation with $a_{1}=a_{2}$ and $b_{1}=b_{2}$ would satisfy (13). Naturally, identification is not possible in this case, which leads to our first condition.
Non-systematic missing pattern: $q_{1}\neq q_{2}$.
Condition (C1) precludes the special case of down-sampling, which is notoriously hard to be treated as a missing data problem.
However, (C1) is inadequate for identification because the equations in (13) are not linearly independent. The lack of uniqueness of the solutions to (13) is at the crux of the identification problem in imputation. The blue lines in Figure 1 plots the set of solutions to (13). Each point on the blue line represents a legitimate TWI imputation whose marginal distributions before and after time $n_{1}$ are the same. Nonetheless, (12) hints an additional requirement for the correct imputation, giving rise to the following condition.
@@
Intuitively, (C2) requires that, given the same neighboring observations, one should impute the missing entries in the same way (at least probabilistically) before and after $n_{1}$. Together, (C1) and (C2) imply a unique solution to the system (13) and are sufficient for identification.
Figure 1: Set of calibrations that correspond to temporal Wasserstein imputation solutions (blue line) and the correct calibrations $(a_{1}^{*},a_{2}^{*})$ and $(b_{1}^{*},b_{2}^{*})$ (red dots). The dashed orange line is the 45 degree line. $q_{1}=3$ and $q_{2}=4$ are used to plot this figure.
@@
Under (C1) and (C2), the only solution to (13) is
$a_{1}=a_{2}=a_{1}^{*},\quad b_{1}=b_{2}=b_{1}^{*}.$
That is, the only 2-dimensional marginal distribution that achieves zero Wasserstein loss and satisfies (C2) is that of the underlying distribution.
As shown in Figure 1, the 45 degree line (dashed orange), which imposes (C2), intersects with the solution set of temporal Wasserstein imputations at the correct calibrations. This example shows, under two natural requirements, TWI can achieve distributional identification through minimizing the Wasserstein distance between the marginal distributions.
In practice, appropriate tuning of $n_{1}$ and $p$ is needed for TWI to find an imputation with marginal distributions approximating the underlying ones. As shown in Proposition 5, it is not necessary to use an $n_{1}$ such that all missing data lie on the same side. Instead, we can select $n_{1}$ from a pre-specified interval $[t_{1},t_{2}]$ which yields the smallest Wasserstein loss, where $t_{2}-t_{1}$, $t_{2}$, and $n-t_{1}$ are sufficiently large. Alternatively, a simple procedure to promote interpolator stability, as suggested by (C2), is to run TWI with different cut-off $n_{1}$, using previously obtained imputations as initializations. This procedure, which we call $k$-TWI, is summarized in Algorithm 2. Finally, as a nonparametric method, $p$ can be chosen to diverge with the sample size with $p/n\rightarrow 0$.
@@ 5 Simulation Studies
In this section, we apply TWI to data generated from various time series models, including multivariate, nonlinear, and nonstationary ones, to examine its performance. For comparison, we employ imputation methods such as Kalman smoothing, the scalar filtering algorithm of Peña and Tsay, (2021), and deterministic methods such as linear and spline interpolations. In addition, the following two missing patterns are considered.
@@
Pattern I simulates sporadic missing data while Pattern II generates patches of missing data. The total length of the time series is set to $n=1000$. Thus, for both missing patterns, 30% of the data are missing. For each data generating process and missing pattern considered, 1000 Monte Carlo simulations are carried out.
We employ the proximal gradient descent (Parikh et al.,, 2014) to efficiently solve step (b) in Algorithm 1 when no side information is available since the projection step is straightforward. Conversely, when $\mathcal{C}$ is defined by a system of linear equations, we solve step (b) using the closed form (6). Throughout we fix $n_{1}=0.4n$ and $p=6$ for TWI. In addition, we apply the $k$-TWI (Algorithm 2) with three cut-off points ($n_{1}\in\{0.25n,0.5n,0.75n\}$) to the simulated data. For the initializations of TWI, both linear interpolation and Kalman smoothing are used in our experiments. Kalman smoothing and the deterministic smoothing techniques are implemented via the imputeTS package (Moritz and Bartz-Beielstein,, 2017) in R. The scalar filter algorithm (denoted as ScalarF henceforth) utilizes intervention analysis, and can be applied to multivariate and nonstationary series. Its implementation details are documented in Section S2 of the supplementary material.

### v3

@@ 4 Theoretical properties
This section investigates the optimization and statistical properties of TWI. We first show Algorithm 1 converges to critical points. We then establish the asymptotic consistency of TWI when the missing data concentrate on one side of $n_{1}$. For general non-systematic missing patterns, a case study of a two-state Markov chain example shows that TWI can identify the underlying marginal distributions under some natural assumptions.
@@ 4.1 Optimization and statistical guarantees
We begin by studying the optimization properties of Algorithm 1. All proofs are relegated to Section S1 of the supplementary material. The following result shows Algorithm 1 is guaranteed to find a local minimum.
@@
[…]
@@
    If, in addition, the cost function $L$ takes the form $L(\mathbf{u},\mathbf{v})=\sum_{k=1}^{p}\ell(u_{k}-v_{k})$, where $\mathbf{u}=(u_{1},\ldots,u_{p})^{\top}$, $\mathbf{v}=(v_{1},\ldots,v_{p})^{\top}$, and $\ell$ is twice differentiable and convex, then $F(\mathbf{w},\mathbf{\Pi})$ is biconvex in $\mathbf{w}$ and $\mathbf{\Pi}$.
Due to the biconvexity of $F$, one can employ off-the-shelf solvers for the subproblems in Algorithm 1. Moreover, it follows directly from the proof that if $\lambda>0$, $F(\mathbf{w},\mathbf{\Pi})$ is $\lambda$-strongly convex in $\mathbf{w}$. Hence, step (b) in Algorithm 1 admits a unique solution in each iteration and can be solved efficiently. It is straightforward to generalize Proposition 3 for multivariate time series, as discussed in Section 3.
Unlike typical non-convex optimization problems, we are only interested in the subset of the global minimizers of $F$ which have the correct marginal distributions (see Section 4.2). The problem simplifies when the missing data happen to lie on the same side of $t=n_{1}$, in which case the global minimizer of $F$ achieves asymptotic consistency.
@@
[…]
@@
Corollary 5.
Assume that the conditions of Proposition 4 hold. Then for continuous $\varphi:\mathbb{R}^{p}\rightarrow\mathbb{R}$ with $|\varphi(x)|\leq C(1+\|x\|^{k})$ for all $x\in\mathbb{R}^{p}$ and some $C<\infty$, we have $\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\varphi(\hat{w}_{t},\ldots,\hat{w}_{t-p+1})\rightarrow\mathbb{E}\varphi(\tilde{x}_{t},\ldots,\tilde{x}_{t-p+1})$ a.s., where $\hat{\mathbf{w}}=(\hat{w}_{0},\ldots,\hat{w}_{n-1})$ is defined in Proposition 4.
@@ 4.2 General missing pattern: A two-state Markov chain example
In this subsection, we study the identification of the underlying marginal distributions via a two-state Markov chain example under the general missing pattern where missing data are not confined to one side of time index $n_{1}$. We connect the non-identification problem to the lack of uniqueness of the solutions to a system of linear equations. We also show that, under two natural conditions, the correct marginal distributions can still be pinned down by TWI. To free our discussion from finite-sample complications, we will focus on the distribution level in this subsection, or intuitively speaking, assume “infinite sample.”
@@
Suppose the time series $\{\tilde{x}_{t}\}_{-\infty<t<\infty}$ under study is a two-state Markov chain. Specifically, $\tilde{x}_{t}\in\{0,1\}$ and the transition probabilities are given by $\mathbb{P}(\tilde{x}_{t}=1|\tilde{x}_{t-1}=0)=p$, and $\mathbb{P}(\tilde{x}_{t}=0|\tilde{x}_{t-1}=1)=q$, where $p,q\in(0,1)$. The associated stationary distribution is denoted by $\lambda_{1}:=\mathbb{P}(\tilde{x}_{t}=0)$ and $\lambda_{2}:=\mathbb{P}(\tilde{x}_{t}=1)$. Consequently, the 2-dimensional marginal distribution of $\{\tilde{x}_{t}\}$ is
$\begin{array}[]{llll}\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(1,1))=\lambda_{2}(1-q),&\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(1,0))=\lambda_{1}p,\\ \mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(0,1))=\lambda_{2}q,&\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(0,0))=\lambda_{1}(1-p).\end{array}$ (9)
@@
We consider the following missing pattern. Without loss of generality, set $n_{1}=0$. Suppose that before time $n_{1}$, there is a missing value every $k_{1}$ observations. Likewise, there is a missing value every $k_{2}$ observations after time $n_{1}$. Specifically, $x_{t}=\bigstar$ if t-t is a positive integer multiple of $k_{1}$, or if $t$ is a positive integer multiple of $k_{2}$, where $k_{1}$ and $k_{2}$ are some integers larger than 2 and $k_{1}\neq k_{2}$ (see condition (C1) below). Next, we parametrize an imputation $\{\hat{w}_{t}\}_{-\infty<t<\infty}$ as follows. Specifically, let
$\mathbb{P}(\hat{w}_{t}=1|(x_{t},x_{t-1})=(\bigstar,1))=a_{1},$ $\mathbb{P}(\hat{w}_{t}=1|(x_{t},x_{t-1})=(\bigstar,0))=b_{1},$ $\mbox{ if }t<n_{1},$ (10)
$\mathbb{P}(\hat{w}_{t}=1|(x_{t},x_{t-1})=(\bigstar,1))=a_{2},$ $\mathbb{P}(\hat{w}_{t}=1|(x_{t},x_{t-1})=(\bigstar,0))=b_{2},$ $\mbox{ if }t\geq n_{1},$ (11)
@@
for some $a_{1},a_{2},b_{1},b_{2}\in[0,1]$. Essentially, we parametrize all possible imputations by $a_{1},a_{2},b_{1},b_{2}$, each corresponding to different 2-dimensional marginal distributions. The correct imputation should calibrate $a_{i}$’s and $b_{i}$’s such that the resulting marginal distributions equal those of the underlying series, given in (9). It is not difficult to see that the correct imputation must choose
$a_{1}^{*}=a_{2}^{*}=1-q,\quad b_{1}^{*}=b_{2}^{*}=p.$ (12)
@@
Note that this calibration does not depend on the missing pattern $k_{1},k_{2}$.
@@
Assuming the optimization is carried out perfectly [^1] This is where we ignore finite sample complications and assume a zero Wasserstein loss is achieved. One may use, for example, the hamming distance as the ground metric for the optimal transport cost., TWI calibrates $a_{i},b_{i},i=1,2$, by equating the 2-dimensional marginal distributions of $\{\hat{w}_{t}\}_{-\infty<t<0}$ and $\{\hat{w}_{t}\}_{0\geq t<\infty}$. Hence, $\{a_{1},a_{2},b_{1},b_{2}\}$ satisfies
$\left.\begin{aligned} \frac{1}{k_{1}}a_{1}-\frac{1}{k_{2}}a_{2}=&(1-q)(\frac{1}{k_{1}}-\frac{1}{k_{2}}),\\ \frac{1}{k_{1}}b_{1}-\frac{1}{k_{2}}b_{2}=&p(\frac{1}{k_{1}}-\frac{1}{k_{2}}).\end{aligned}\right\}$ (13)
Derivation of (13) is relegated to the Section S1 of the supplementary material. Now we introduce two conditions.
Non-systematic missing pattern: $k_{1}\neq k_{2}$.
@@
Clearly, identification is not possible when $k_{1}=k_{2}$, in which case any imputation with $a_{1}=a_{2}$ and $b_{1}=b_{2}$ would satisfy (13), so (C1) is necessary for identification. Intuitively, (C1) precludes the special case of down-sampling, which is notoriously hard to be treated as a missing data problem. However, even with (C1) the solutions to (13) are not unique. This lack of uniqueness is at the crux of the identification problem for time series in the distribution matching framework (Muzellec et al.,, 2020; Wang et al.,, 2025). There are many imputations whose marginal distributions before and after $n_{1}$ are the same, but they may not equal to the one given in (9). Inspired by (12), (C2) imposes an additional requirement for the ideal imputation, which can be verified without knowledge about the missing data. It requires that, given the same neighboring observations, one should impute the missing entries in the same way (at least probabilistically) before and after $n_{1}$. Together, (C1) and (C2) imply a unique solution to the system (13), which guarantees identification.
@@
Under (C1) and (C2), the only solution to (13) is $a_{1}=a_{2}=a_{1}^{*}$, $b_{1}=b_{2}=b_{1}^{*}$. That is, the only 2-dimensional marginal distribution that achieves zero Wasserstein loss and satisfies (C2) is that of the underlying distribution.
As shown in Proposition 6, it is not necessary to use an $n_{1}$ such that all missing data lie on the same side. In practice, appropriate tuning of $n_{1}$ and $p$ is needed. We can select $n_{1}$ from a pre-specified interval $t_{1},t_{2}]$ which yields the smallest Wasserstein loss, where $t_{1}$ and $n-t_{2}$ are sufficiently large. Alternatively, a simple procedure to promote interpolator stability is to run TWI with different splitting cut-offs, using previously obtained imputations as initializations. This procedure, which we call $k$-TWI, is summarized in Algorithm 2. Finally, similar to other nonparametric methods, $p$ can be chosen to diverge with the sample size so long as $p/n\rightarrow 0$.
@@ 5 Simulation Studies
In this section, we apply TWI to data generated from various time series models, including multivariate, nonlinear, and nonstationary ones, to examine its finite sample performance. For comparison, we employ Kalman smoothing, the iterative gap filling by SSA (Kondrashov and Ghil,, 2006), the scalar filtering algorithm (Peña and Tsay,, 2021), and linear interpolations. The following two missing patterns are considered:
@@
Pattern I simulates data missing at random while Pattern II generates patches of missing data within the same positions of every 20 observations. The total length of the time series is set to $n=1000$. Thus, for both missing patterns, 30% of the data are missing. For each model and missing pattern considered, 1000 Monte Carlo simulations are carried out.
The proximal gradient descent (Parikh et al.,, 2014) is used to efficiently solve step (b) in Algorithm 1 when no side information is available because the projection step is straightforward. When $\mathcal{C}$ is defined by a system of linear equations, we solve step (b) using the closed form (6). The POT package (Flamary et al.,, 2021) in Python is used to solve the discrete optimal transport problem. Throughout we fix $n_{1}=0.4n$ and $p=6$ for TWI. In addition, we apply the $k$-TWI (Algorithm 2) with three cut-off points $n_{1}\in\{0.25n,0.5n,0.75n\}$. For the initializations of TWI, both linear interpolation and Kalman smoothing are used in our experiments. Kalman smoothing and gap filling with iterative SSA (iSSA henceforth) are implemented, respectively, via the imputeTS (Moritz and Bartz-Beielstein,, 2017) and Rssa (Golyandina et al.,, 2018) packages in R. As discussed in Section 3.3, the window length $L$ for which iSSA can be implemented is restricted. Here we use $L=6$ for iSSA. In the reconstruction stage for iSSA, the first three eigentriples are used for grouping with the linear interpolation as initialization. The scalar filter algorithm (denoted as ScalarF henceforth) utilizes intervention analysis, which is detailed in Section S2 of the supplementary material. Codes for the numerical experiments in this and the next section can be found in the author’s Github page: https://github.com/shuochieh/Wasserstein_imputation.